# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

### v1

@@ 7 Summary and future work
Fig. 19: Parameter convergence, clean data, 151 data points, ensemble: Convergence of the four unknown parameters of the engine showing the range of variation across the 30 independent runs for (a) Baseline PINN (hybrid PINN) model ($t=0-1$ min), (b) Multi-stage transfer learning model ($t=3-4$ min), and (c) Few-shot transfer learning model ($t=3-4$ min) for clean data case with 151 data points in the loss function. The min-max band indicates the range over which the parameter estimation varies across the thirty independent runs that we used in this study. We observed that this variation is reduced significantly as training progresses, indicating a good repeatability of our PINN framework.
Input data and data generations
@@ Acknowledgement
Conflict of interest
Author George Em Karniadakis has received research support from Cummins Inc. Karniadakis has financial interests with the company PredictiveIQ. The rest of the authors declare no competing interests.
@@ Appendix A Engine model
For the value considered for the parameters in the equation and general workflow, interested readers may refer to our previous work [20].
Appendix B Additional figures and Table
B.1 Multi-stage transfer learning prediction results for clean data
We provide additional results showing the predictions for different engine related QOIs. The plots combine the results from 30 independent runs and represent an estimate of the epistemic uncertainty of the network.
Fig. B.1.1: Violin plot, multi-stage transfer learning, clean data, 301 data points showing the variation in engine parameter predictions over 30 runs with clean data and 301 data points. More data points help reduce the variation in estimation across all data segments as compared to the model trained with 151 data points.
Fig. B.1.2: Predicted dynamic states, multi-stage transfer learning, clean data, 151 data points, ensemble: Ensemble prediction results for primary gas flow dynamics states from transfer learning model (data scale normalized due to confidentiality). The plot shown is for scenario where clean data is used with 151 data points used for online training. The ensemble mean (red) for the 30 runs is plotted against ground truth (black) for the eight dynamic states defined by ODE’s describing the gas flow dynamics. Error bands (shown in light green) indicate the spread of data between point-wise maximum and minimum values across all 30 runs.
Fig. B.1.3: Predicted gas flow, multi-stage transfer learning, clean data, 151 data points, ensemble: Ensemble prediction results for gas flow rates from the Transfer learning model with clean data over 30 independent runs across all segments. 151 data points were used during online parameter estimation task.
Fig. B.1.4: Predicted empirical states, multi-stage transfer learning, clean data, 151 data points, ensemble: Ensemble prediction estimates for empirical relationships for clean data with Transfer learning model. Note that in region $\approx 1.5-3.5\,min$, the model’s prediction indicates a spectral bias in learning the high frequency signals that typically result from engine idling. Elsewhere, the model predictions show good match with target values for these empirical relationships.
B.2 Multi-stage transfer learning prediction results for noisy data
Fig. B.2.1: Predicted dynamic states, multi-stage transfer learning, noisy data, 301 data points, ensemble: Ensemble prediction results for primary gas flow dynamics states from Transfer learning model (data scale normalized due to confidentiality). The plot shown is for scenario where noisy data is used with 301 data points used for online training. The ensemble mean (red) for the 30 runs is plotted against ground truth (black) for the eight dynamic states defined by ODE’s describing the gas flow dynamics. Error bands (shown in light green) indicate the spread of data between point-wise maximum and minimum values across all the 30 runs.
Fig. B.2.2: Predicted gas flow rate, multi-stage transfer learning, noisy data, 301 data points, ensemble: Ensemble prediction results for gas flow rates from the Transfer learning model with noisy data. 301 data points were used during online parameter estimation task.
Fig. B.2.3: Predicted empirical states, multi-stage transfer learning, noisy data, 301 data points, ensemble: Ensemble prediction estimates for empirical relationships for noisy data with transfer learning model with 301 data points. Note that in region $\approx 1.5-3.5\,min$, the model’s prediction indicate a spectral bias in learning the high frequency signals that typically result from engine idling. Also, there is a significant increase in expected variability in prediction of these empirical states under noisy conditions as compared to clean conditions shown in figure B.1.4, correlating with increase in variability in prediction of other states and parameters as well.
B.3 Few shot transfer learning for clean data
Fig. B.3.1: Violin plot, few-shot learning, clean data, 301 data points: of the predicted value of the unknown parameters. 301 data points are considered for each of the known parameters (for each one minute segment) and no noise is considered here. The violin plots are clipped at the maximum and minimum value of the predicted parameters. The dots over the violin plots show the swarm plot indicate predicted value for individual runs. The grey region shown between the dotted lines indicates the $\pm$ 7.5% error band. The middle dotted line indicates the true value of the unknown parameter, the top dotted line indicates 1.075 times the true value while the bottom dotted line indicates 0.925 times the true value.
Fig. B.3.2: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for state variable and $x_{r}$ and $T_{1}$ for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.3.3: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for different variables for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.3.4: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for the empirical formulas for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
B.4 Few shot transfer learning for noisy data
Fig. B.4.1: Predicted states, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for state variable and $x_{r}$ and $T_{1}$ for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.4.2: Predicted states, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for different variables for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. B.4.3: Predicted states, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for the empirical formulas for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Appendix C Additional Tables

### v2

@@ 5 Comparison between Multi-stage transfer learning and Few-shot transfer learning
Fig. 14: Parameter convergence, clean data, 151 data points, ensemble: Convergence of the four unknown parameters of the engine showing the range of variation across the 30 independent runs for (a) Baseline PINN (hybrid PINN) model (t=01t=0-1 min), (b) Multi-stage transfer learning model (t=34t=3-4 min), and (c) Few-shot transfer learning model (t=34t=3-4 min) for clean data case with 151 data points in the loss function. The min-max band indicates the range over which the parameter estimation varies across the thirty independent runs that we used in this study. We observed that this variation is reduced significantly as training progresses, indicating a good repeatability of our PINN framework.
6 Limitations and future work
Despite the demonstrated efficacy and advancements of our hybrid approach, there exist certain limitations of this method that we acknowledge. First, the systemic uncertainty has been modelled using fixed dropout rates in our empirical models that represent outputs from largely independent systems. The dropout rate in this study was chosen heuristically, and future work will explore learning data-driven dropout schedules for these systems. Second, while our few-shot TL strategy reduces fine-tuning costs, its performance is sensitive to the representativeness of the pre-training dataset and requires a copious amount of data for effective learning. Third, we note that the input data obtained from field measurements were collected under normal loading conditions for a range of engine speeds and actuator conditions. The current model may require fine-tuning of hyperparameters (such as self-adaptive weights) to deal with special loading conditions that may be encountered in the field. However, we believe that adapting to new scenarios will likely require marginal effort due to the physics-based regularization embedded in our framework. Lastly, our future research work will explore methods that can help reduce the aleatoric uncertainty leading to high variability in parameter estimation under noisy conditions. While the proposed methods show good accuracy for clean data, we observed a few instances (runs) of overfitting while using noisy data. To control this overfitting, one can consider early stopping and a good learning rate schedule. To further mitigate these issues, complementary robustness strategies can be considered in future work. For example, data augmentation using input perturbations or noise realizations can help improve generalization in our few-shot method. Adversarial training provides another approach to enhance model robustness by optimizing the model against worst-case perturbations. Additionally, robust regularization or loss formulations can also suppress overfitting and improve model’s performance. Together, these methods provide a complementary pathway for improving the reliability and real-world resilience of the proposed few-shot transfer learning framework.
7 Summary
In this study, we developed two computationally efficient frameworks for parameter estimation in diesel engine subsystems by integrating Physics-Informed Neural Networks (PINNs) with neural operator and transfer learning strategies. The proposed approach advances the state of the art in physics-based fault diagnostics and health monitoring of diesel engines by addressing two primary limitations of existing (conventional and neural network-based) methods. Firstly, the proposed method addresses the need to create a holistic framework for predicting engine dynamics at a system level, which encompasses all critical air flow sub-systems simultaneously. Secondly, our framework addresses a key challenge of incorporating models grounded in physical knowledge with measurable sensory data to estimate gas flow dynamics, along with the identification of unknown system parameters for use in continuous health monitoring during system operation.
To evaluate the applicability of this methodology in real scenarios, we introduced systematic variability through dropouts in pre-trained empirical networks for generating stochastic outputs from these networks. Additionally, we presented a hybrid framework that leverages Deep Operator Networks (DeepONet) for accelerated state predictions, with two transfer learning (TL) strategies designed to reduce computational time to meet requirements for field use. The first strategy, a multi-stage TL method relies on knowledge transfer between our PINN-based neural networks when learning system dynamics and parameter estimation across different temporal segments of the input signal. The second strategy, a few-shot TL approach leverages a large pre-trained neural network to generalize estimation of gas dynamics across multiple operating conditions simultaneously. Our results, based on simulated data with field-derived input signals and corrupted with Gaussian noise to mimic aleatoric uncertainty in sensory signals demonstrate that both methods effectively estimate unknown system parameters and dynamics states across diverse operating conditions. Notably, under noisy conditions, the parameter estimates exhibit variability in both methods which is expected since this challenge is shared by most existing approaches. Nonetheless, the proposed framework, demonstrates significant promise towards the use of physics-based methods in engine health monitoring and is a major step forwards as compared to prior works:
  • Compared to pure data-driven methods such as [8, 20, 14], our method offers improved interpretability by estimating physical parameters of engine sub-systems which are well defined by physical models. Compared to existing methods where a residual signal is used for fault estimation, our proposed method identifies scalar parameters from measured sensory data in an inverse problem paradigm. Drifts in these physical parameters can be monitored for triggering a fault as compared to a temporal residual signal, where momentary drifts over a short segment could lead to false positives. Further, existing methods for predicting engine dynamics are primarily focused on monitoring individual sub-components and their behaviour. On the contrary, our hybrid approach provides simultaneous estimates for gas flow dynamics and unknown system parameters for the complete engine model. This is a significant advancement for building comprehensive strategies for fault and health monitoring using system-level information in series rather than analysing individual component blocks for anomalies.
  • Compared to a baseline PINNs implementation [30], our hybrid approach advances it further since it reduces the computational time as independent state variables are approximated using offline trained neural operators, DeepONet in this case. While the hybrid multi-stage TL approach reduces computational time by approximately 50%, the few-shot TL paradigm further improves the computational time by an additional 25% in comparison to the hybrid PINN model.
  • By fusing physics-based knowledge with measured sensory data, our proposed method provides a framework for extending use of neural network-based models across different operating conditions which otherwise is challenging. To test robustness of our framework, we use different sets of unknown parameters to generate data for pre-trained networks and DeepONet while inferring an unseen set of parameters with PINN learning. Existing neural network approaches that learn purely empirical mappings between inputs and outputs will likely produce erroneous estimates when evaluated at unseen conditions.
The key scientific and practical value added by this study lies in the fusion of physics-based neural networks with operator learning (data-driven) and transfer learning strategies as a step towards real-time, explainable, and unambiguous health monitoring for diesel engines. By transitioning from a residual signal-based fault detection paradigm, this framework lays the groundwork for system-wide fault detection and mitigation strategies that are rooted in physics-based knowledge as well as measured sensory data from the engine. Hence, we believe such a framework is well-suited for developing digital twin strategies for modelling diesel engines in the future.
Input data and data generation
@@ Acknowledgement
  • Appendix A: A description of the engine model considered in this study.
  • Appendix B: A brief description of vanilla PINN model
  • Appendix C: Detailed flow chart for the hybrid PINN model
  • Appendix D: Explanation of different loss terms used for training are discussed.
  • Appendix E: Detail and results of DeepONet for approximation of EGR and VGT states.
  • Appendix F: Additional results for multi-stage transfer learning.
  • Appendix G: Additional results for few-shot transfer learning.
@@ Appendix A Engine model
For the value considered for the parameters in the equation interested readers may refer to previous work [30]. We have shown detail flow chart for calculation of physics-informed losses in Appendix C.
Appendix B Introduction to PINN
In the present study, we have considered Physics-Informed Neural Networks (PINNs). In this section, we briefly discuss the basic idea of PINN. PINN consists of two main parts: an approximation of a solution of the state variable of a PDE/ODE using a neural network, and the second part is the calculation of a physics-informed loss function. We will discuss it in terms of ODE, as in the present study, our problem is confined to ODEs only, and the concept can be translated to PDEs. PINNs have two different paradigms, the first one is the solution of a differential equation with known initial (and boundary) conditions, which we call the forward problem. The second one is the evaluation of unknown parameters ($\bm{\Lambda}$) (field) from measured data, which is known as an inverse problem. The neural network approximate the state variable $y$ whose input is time $t$. Thus,
$y=\mathcal{N}_{y}(\bm{\theta},t)$ (A.51)
where $\mathcal{N}_{y}$ represent the neural network parametrized with $\bm{\theta}=\{\bm{W},\bm{b}\}$ are the weights ($\bm{W}$) and biases ($\bm{b}$) of the neural network. The derivative of the state variable, $\dfrac{dy}{dt}$ is generally calculated using automatic differentiation [40] of $y$ with respect to input $t$. Once the derivative of the state variable is calculated, the next step is to calculate the residual of the differential equation. This is done by adding the derivative and other parts in the differential equation. A PINN model has two kinds of loss functions. The first one is the Physics-informed loss, which is a function of the residual (generally MSE loss of the residue). The second one is the data loss. In the case of the forward problem, the data loss is due to initial condition (and boundary condition, in case of PDEs), which is generally the MSE loss of the data. In case of the inverse problem, there is an additional data loss due to the known value of the state (or derived) variable. The total loss is given as the weighted sum of the physic loss and data loss,
$\mathcal{L}(\bm{\theta})=\lambda_{1}\mathcal{L}_{\text{data}}+\lambda_{1}\mathcal{L}_{\text{phy}}$ (A.52)
where $\lambda_{1}$ and $\lambda_{2}$ are weights which can be fixed or adaptive [36]. The parameters $\bm{\theta}$ in the case of a forward problem or $\bm{\theta}$ and $\bm{\Lambda}\}$ in the case of an inverse problem and the weights ($\lambda$) may be optimized using a min-max problem.
$\min\limits_{\bm{\theta},\bm{\varLambda}}\;\max\limits_{\lambda}\mathcal{L}(\bm{\theta},\varLambda)$ (A.53)
$\bm{\theta},\bm{\varLambda}$ $=\bm{\theta},\bm{\varLambda}-lr_{\theta,\varLambda}\nabla_{\theta,\varLambda}\mathcal{L}(\bm{\theta},\varLambda)$ (A.54a)
$\lambda$ $=\lambda+lr_{\lambda}\nabla_{\lambda}\mathcal{L}(\bm{\theta},\varLambda),$ (A.54b)
where $lr_{\theta}$ denotes the network optimiser learning rate for $\bm{\theta}$, $lr_{\varLambda}$ represents the learning rate for the optimizer used to identify unknown scalar parameters $\bm{\varLambda}$, and $lr_{\lambda}$ refers to the learning rate for the optimizer used for updating self-adaptive weights. Incorporating distinct learning rates and optimizers allows for fine tuning of each component for efficient convergence.
Appendix C Detailed flow chart
We have discussed the problem setup in section 1.2, and in subsequent sections, we have discussed how we can approximate different variables using DNNs. Further, we discussed the basic philosophy and implementation of PINN along with DeepONet in the present problem context. We present a flowchart for calculating the physics-informed loss for the present problem in Fig. A-1. We also note that in the flowchart, we have not shown the data loss and self-adaptive weights. We briefly summarize below the sequence of training for various parts of the proposed workflow,
  1. 1.
    DeepONet training: We consider two DeepONets, the first one predicts the two EGR states $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$, and the second DeepONet predicts the VGT state $\tilde{u}_{vgt}$. These are trained offline using laboratory data as discussed in Section 2.3 and Appendix E. The DeepONets are considered pre-trained in the PINN formulation, and the parameters of the DeepONets are fixed during PINN training.
  2. 2.
    Training of DNN for empirical formulae: We consider six DNN to approximate the empirical formula for $\eta_{vol}$, $f_{egr}$, $F_{vgt,\Pi_{t}}$, $\eta_{tm}$, $\eta_{c}$ and $\Phi_{c}$ as discussed in Section 2.2. These DNNs are trained offline using laboratory data. Similar to the DeepONets, the parameters of these DNNs are also fixed during the PINN training; however, the dropout is considered.
  3. 3.
    PINN training for inverse problem: We consider four networks to approximate the state variables $p_{em}$, $p_{im}$, $\omega_{t}$ and $x_{r}$ and $T_{1}$. We formulate the hybrid PINN model as discussed in section 2.4, combining the pre-trained DeepONets for the EGR and VGT states and the pre-trained DNNs for the empirical formulae. In the hybrid PINN, the trainable parameters are parameters of the PINN network, unknown parameters for the inverse problem, and the self-adaptive weights. The parameters of the DeepONet and DNN for the empirical formula are kept fixed in the hybrid PINN.
We consider two different kinds of datasets, the laboratory data and the field data. The laboratory data are used to train the pre-trained networks, i.e., the DeepONet and the DNN for empirical formulas. The field data are used in the inverse problem using PINN. We also consider two transfer learning paradigms,
  • Multi-stage TL: In this approach the hybrid PINN model is trained for the first segment of time ($0\sim 60$ sec). In the subsequent time segments, we considered a multi-stage transfer learning as discussed in Section 2.5.
  • Few-shot TL: Few-shot TL consists of two phases. In the first phase, we trained the multi-head networks for state variables $p_{em}$, $p_{im}$, $\omega_{t}$ and $x_{r}$ and $T_{1}$ offline using data loss function. In this phase, the multi-head networks are trained using laboratory data. In the second phase, we replace the head and train only the head (output layer) of these networks using hybrid PINN as discussed in Section 2.6. We also like to mention that in Few-shot TL we calculate the derivative of the output only once and keep it outside the training loop as discussed in Section 2.6. In the second phase, the PINN networks are trained using field data. The trainable parameters are the weights and biases of the output layers of the PINN network, the unknown parameters, and the self-adaptive weights.
Fig. A-1: Flow chart for proposed hybrid-PINNs model: Flow chart for the proposed hybrid PINN model for the inverse problem for the engine for prediction of dynamics of the system variables and estimation of unknown parameters. The inputs are input control vector $\{u_{\delta},u_{egr},u_{vgt}\}$ and engine speed $n_{e}$. Four neural network $\mathcal{N}_{i}(t;\bm{\theta})$, $i=1,2,3,4$ indicated in dashed rectangular oval takes time $t$ as input and predict $p_{im}$, $p_{em}$, $x_{r}$ and $T_{1}$ and $\omega_{t}$. Two DeepONets $\mathcal{D}_{i}(.;\bm{\theta}_{DPi}$, $i=1,2$ indicated in double line rectangular pentagon takes time $t$ and others as input and predict $\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$ and $\tilde{u}_{vgt}$ as shown in Table 3. The parameters (weights and biases) of these pre-trained DeepONets are kept fixed to predict the three states. Four unknown parameters indicated in octagon are $\eta_{sc}$, $h_{tot}$, $A_{egrmax}$ and $A_{vgtmax}$. Six pre-trained neural networks $\mathcal{N}_{i}^{(P)}(.;\bm{\theta})$, $i=1,2,\dots,6$ indicated in dashed-dotted rectangular oval takes appropriate input and predict the empirical formulae as shown in Table 2. The parameters (weights and biases) of these pre-trained DNNs are kept fixed to predict the empirical formulae. There are eight main blocks calculating different variables. The equations for the calculation of each of the quantities are shown in Appendix A. Cylinder flow: calculates $W_{ei}$, $W_{f}$ and $W_{eo}$ using Eqs. A.3, A.5 and A.4 respectively. Cylinder Temperature: calculates $x_{v}$, $x_{p}$, $T_{e}$ and $T_{em}$ using Eqs. A.12, A.11, A.7 and A.15 respectively. $h_{tot}$ and $\eta_{sc}$ are considered as learnable parameters in the calculation of $T_{em}$ and $T_{e}$ respectively. EGR dynamics: calculated $\tilde{u}_{egr}$ using Eq. A.18. EGR Flow: calculates EGR mass flow $W_{egr}$ using Eq. A.19. $A_{egrmax}$ is considered as learnable parameter. Compressor flow: calculates compressor mass flow $W_{c}$ using Eq. A.49. Compressor Power: calculates compressor power $P_{c}$ using Eq. A.39.Turbine Flow: calculates turbine mass flow $W_{t}$ using Eq. A.26. $A_{vgtmax}$ is considered as trainable parameter. textbfTurbine Power: calculates effective turbine power $P_{t}\eta_{m}$ using Eq. A.32 There are three blocks, which calculate the residual of the equation. • The first block calculates the residual for state equations for $p_{im}$ and $p_{em}$, • The second one calculates the residual for the equations of $x_{r}$ and $T_{1}$, • The third block calculates the residual for the state equation for $\omega_{t}$. There is another block that calculates the physics losses. The calculation for data losses are not shown. The figure is adopted from [30].
Appendix D Detailed loss function for engine model
In section 1.2, we discussed the inverse problem of interest in this study. Our objectives are to: (1) learn the dynamics of the six gas flow dynamics state estimates, three of which are approximated using two DeepONets, and (2) infer the unknown parameters in the system, given field measurements of system variables $\{p_{im},p_{em},\omega_{t},W_{egr},T_{em}\}$ as well as Eqs. (3)-(10), using PINNs. To achieve these objectives, we utilize four DNNs as surrogates for the solutions to different equations, as shown in Table 4. We encode physics loss for Eqs (3)-(10) (EGR, VGT actuator equations are not included as these are approximated using DeepONets) using the automatic differentiation and formulate the physics-informed loss. The total loss function considered for training the PINNs model is given by:
$\begin{split}\mathcal{L}(\bm{\theta},\bm{\mathcal{\varLambda}},\bm{\lambda}_{p_{im}},\bm{\lambda}_{p_{em}},\bm{\lambda}_{\omega_{t}},\bm{\lambda}_{W_{egr}},\lambda_{T_{em}},\lambda_{T_{1}})=&\mathcal{L}_{p_{im}}+\mathcal{L}_{p_{em}}+\mathcal{L}_{\omega_{t}}+100\times\mathcal{L}_{x_{r}}+\lambda_{T_{1}}\times\mathcal{L}_{T_{1}}+\\ &\mathcal{L}^{ini}_{p_{im}}+\mathcal{L}^{ini}_{p_{em}}+\mathcal{L}^{ini}_{\omega_{t}}+100\times\mathcal{L}^{ini}_{x_{r}}+100\times\mathcal{L}^{ini}_{T_{1}}+\\ &\mathcal{L}^{data}_{p_{im}}(\bm{\lambda}_{p_{im}})+\mathcal{L}^{data}_{p_{em}}(\bm{\lambda}_{p_{em}})+\mathcal{L}^{data}_{\omega_{t}}(\bm{\lambda}_{\omega_{t}})+\\ &\mathcal{L}^{data}_{W_{egr}}(\bm{\lambda}_{W_{egr}})+\lambda_{Tm}\mathcal{L}^{data}_{T_{em}},\end{split}$ (A.55)
The individual physics losses are represented as:
$\mathcal{L}_{p_{im}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(p_{im}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{dp_{im}^{(k)}}{dt}-\dfrac{R_{a}T_{im}}{V_{im}}\left(W_{c}^{(k)}+W_{egr}^{(k)}-W_{ei}^{(k)}\right)\right)^{2}$ (A.56a)
$\mathcal{L}_{p_{em}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(p_{em}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{dp_{em}^{(k)}}{dt}-\dfrac{R_{e}T_{em}^{(k)}}{V_{em}}\left(W_{eo}^{(k)}-W_{t}^{(k)}-W_{egr}^{(k)}\right)\right)^{2}$ (A.56b)
$\mathcal{L}_{\omega_{t}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(\omega_{t}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(\dfrac{d\omega_{t}^{(k)}}{dt}-\dfrac{P_{t}^{(k)}\eta_{m}-P_{c}^{(k)}}{J_{t}\omega_{t}^{(k)}}\right)^{2}$ (A.56c)
$\mathcal{L}_{x_{r}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(x_{r}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(x_{r}^{(k)}-\dfrac{(\Pi_{e}^{(k)})^{1/\gamma_{a}}(x_{p}^{(k)})^{-1/\gamma_{a}}}{r_{c}x_{v}^{(k)}}\right)^{2}$ (A.56d)
$\mathcal{L}_{T_{1}}=$ $\dfrac{1}{n}\sum_{k=1}^{n}r(T_{1}^{(k)})^{2}=\dfrac{1}{n}\sum_{k=1}^{n}\left(T_{1}^{(k)}-\left(x_{r}^{(k)}T_{e}^{(k)}+(1-x_{r}^{(k)})T_{im}\right)\right)^{2}$ (A.56e)
$\mathcal{L}^{data}_{\phi}(\bm{\lambda}_{\phi})=$ $\dfrac{1}{n_{\phi}}\sum_{j=1}^{n_{\phi}}\left[\left(\phi_{data}^{(j)}-\hat{\phi}_{data}^{(j)}\right)\lambda_{\phi}^{(j)}\right]^{2},\quad\quad\phi=p_{im},p_{em},\omega_{t},W_{egr}$ (A.57a)
$\mathcal{L}^{data}_{T_{em}}=$ $\dfrac{1}{n_{T_{em}}}\sum_{j=1}^{n_{T_{em}}}\left[\left(T_{em_{data}}^{(j)}-\hat{T}_{em_{data}}^{(j)}\right)\right]^{2}$ (A.57b)
The initial condition losses are given as:
$\mathcal{L}^{ini}_{\phi}=\dfrac{1}{1}\sum_{j=1}^{1}\left(\phi_{0}^{(j)}-\hat{\phi}_{0}^{(j)}\right)^{2},\;\;\;\;\;\;\;\;\phi=p_{im},p_{em},\omega_{t},x_{r},T_{1}$ (A.58)
Appendix E DeepONet: Approximation of EGR states ($\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$) and VGT states ($\tilde{u}_{vgt}$) using DeepONets
In section 2.3, we discussed how the states describing the EGR actuator ($\tilde{u}{egr1}$ and $\tilde{u}{egr2}$) and the VGT actuator ($\tilde{u}{vgt}$) dynamics are approximated using two DeepONets. Fig. A-2 illustrates a schematic diagram of the DeepONet architecture for the EGR actuator states. In this configuration, the DeepONet comprises of one trunk, which takes time as input, and four branch networks. The four branch networks receive $u_{d}=u_{degr}=u_{egr}(t-\tau_{degr})$, $\tilde{u}_{egr1,0}$, $u_{d}=u_{degr}=u_{egr}(t-\tau_{degr})$ and $\tilde{u}_{egr2,0}$ as inputs, respectively. For the VGT actuator approximation, there are only two branch networks that take $u_{d}=u_{dvgt}=u_{vgt}(t-\tau_{dvgt})$ and $\tilde{u}_{vgt,0}$ as input. As discussed earlier, we have considered Causality-DeepONet [39] for these DeepONets. The input-output data structures for the DeepONet model are shown in Table A-1.
$\mathcal{L}_{i}(\bm{\theta}_{DP,i})=$ $\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\hat{y}_{i}^{(j)}\right]^{2},\;\;\;\;\;i=1,2$ (A.59a)
$=$ $\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\mathcal{D}_{i}(t,\bm{x}_{i};\bm{\theta}_{DP,i})^{(j)}\right]^{2},\;\;\;\bm{x}\rightarrow\text{Branch inputs}$ (A.59b)
We also note that the DeepONet for the EGR states has two output quantities. Consequently, the loss function for this DeepONet is the weighted summation of the loss function for individual output quantities. These networks are trained using Adam optimizer. As discussed earlier, we utilize simulation data to train the DeepONets and subsequently replace the EGR and VGT dynamics with these pre-trained networks within the PINN model. Note that for DeepONet training here, we do not incorporate dropouts since the outputs of these DeepONets are used as inputs to the pre-trained empirical networks that include dropouts to inherently addresses the variability, as discussed in section 2.2. In the following section, we present details about our hybrid PINN framework to tackle the parameter estimation problem.
The labelled data considered for the training and testing of these DeepONet are listed in Table 5. We divide the problem into a prediction of response of one minute. Thus, the training dataset consists of data of “Case-I” divided into 120 samples of ${0-60},{60-120},\dots,$ sec and “Case-VI”, divided into 24 samples of ${0-60},{60-120},\dots,$ sec. We also consider data augmentation in the training dataset by considering a sliding window of the starting point of the time series. The testing data set “Case-V” is also divided into 20 samples of ${0-60},{60-120},\dots,$ sec. By doing this, we convert the problem to an initial value problem of one-minute duration. We consider the input signal $u_{egr}$ to be recorded at equal $dt=0.2$ sec. The output $\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$ are also recorded at same $dt=0.2$ sec. The trunk takes time input $t=0.0,\;0.20,\;0.4,\;\dots,\;60.0$ sec. We know the states are bounded; thus we consider output transformation and scaling so that the outputs are always within the limit. The network sizes and other details are shown in Table A-1.
The loss function for individual output variables is the mean square error between the true value ($\tilde{u}_{1}$) and the predicted one ($\tilde{u}_{1}(\bm{\theta})$). The total loss is the weighted sum of losses for both outputs. For EGR actuator dynamics, we use weights $a_{1}=1$ and $a_{2}=100$ (ref. Fig. A-2). The DeepONets are trained using the Adam optimizer, with $3\times 10^{5}$ epochs for DeepONet $\mathcal{D}_{1}$ and $4\times 10^{5}$ epochs for DeepONet $\mathcal{D}_{2}$. Relative $L_{2}$ errors are as shown in Table A-2. The mean and standard deviation of the error are calculated from six independent training sessions and show good accuracy. To further validate the model’s robustness, we added a $3\%$ Gaussian noise to the labelled output data and trained the networks with this noisy data. The relative $L_{2}$ error for these cases is also shown in Table A-2, and we note only a marginal change in error, demonstrating the model’s robustness to noise in training data.
Fig. A-2: Schematic diagram of DeepONet for prediction of EGR states showing branches and trunk. The trunk network takes time as its input, at which the value of the output function needs to be evaluated. The four branches take $u_{d}$, $\tilde{u}_{1,0}$, $u_{d}$ and $\tilde{u}_{2,0}$ as input and approximate two outputs $\tilde{u}_{1}=\tilde{u}_{egr1}$ and $\tilde{u}_{1}=\tilde{u}_{egr2}$. Here, $\tilde{u}_{1,0}=\tilde{u}_{egr1,0}$ and $\tilde{u}_{2,0}=\tilde{u}_{egr2,0}$ are the initial conditions and $u_{d}=u_{degr}=u_{egr}(t-\tau_{degr})$, where $\tau_{degr}$ is the time delay coefficient. The loss function for individual output variables is the mean square error between the true value ($\tilde{u}_{1}$) and the predicted one ($\tilde{u}_{1}(\bm{\theta})$), and the total loss is the weighted sum ($a_{1}=1$, $a_{2}=100$) of loss for both outputs. In the case of prediction of VGT actuator states, there are only two branch networks, branch 1 and 2 and only one output $\tilde{u}_{1}=\tilde{u}_{vgt}$. The inputs to these branch networks are $u_{d}=u_{dvgt}=u_{egr}(t-\tau_{dvgt})$ and the initial condition $\tilde{u}_{1,0}=\tilde{u}_{vgt,0}$. Appropriate scaling of the input and output are considered for fast and better training. Details of the network sizes and input-output data structures are shown in Table A-1.